OpenStudy (anonymous):

A man bought 42 stamps, some 13¢ and some 18¢. How many of each kind did he buy if the cost was $6.66?

OpenStudy (anonymous):

Let 13¢ stamp = x, 18¢ stamp = y x + y = 42 0.13x + 0.18y = 6.66 x = 42 - y 0.13(42 - y) + 0.18y = 6.66 Solve for y, then for x.

OpenStudy (anonymous):

x=18 y=21